# Well-posedness and stability of the Lagrange representation of the n-D wave equation via boundary triples

### v1

@@ 3. Well-posedness
In [15] it was shown that there is a boundary triple associated to the wave equation in the Dirac representation. In this section we show that we can also associate a boundary triple to the wave equation in the Lagrange representation. This will allow us to parameterize all maximally dissipative boundary conditions.
Recall the decomposition of $\partial\Omega$ into $\Gamma_{0}$, $\Gamma_{1}$, $\Gamma_{2}$ $\Gamma_{3}$ and $\Gamma_{4}$. Since the case where $\Gamma_{0}$ is equal to $\partial\Omega$ simply corresponds to the wave equation with Dirichlet boundary, which is well-studied, we exclude that case. In particular, we will see, that the boundary conditions (3) is one of these maximally dissipative boundary conditions. For the following we only have to distinguish between $\Gamma_{0}$ and the rest. Hence, let us abbreviate $\tilde{\Gamma}\coloneqq\partial\Omega\setminus\overline{\Gamma_{0}}$, which is essentially the same as $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\cup\Gamma_{4}$.
@@
$X\coloneqq\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\times\mathrm{L}^{2}(\Omega).$
We use the equivalent norm for $\mathrm{H}^{1}(\Omega)$ given by the Poincaré inequality Theorem C.1 and define
$\langle x,y\rangle_{X}\coloneqq\big{\langle}\tfrac{1}{\rho}x_{2},y_{2}\big{\rangle}_{\mathrm{L}^{2}(\Omega)}+\langle T\nabla x_{1},\nabla y_{1}\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}x_{1},\gamma_{0}y_{1}\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}\text{.}$
Equivalence to the canonical inner product on $X$ is assured by virtue of the Poincaré inequality and the assumptions on $\rho$, $T$ and $k_{1}$. This is the energy inner product that is favored in port-Hamiltonian formulations plus the inner product on the boundary, which is necessary to prevent constant functions in the second argument to have vanishing norm. We arrive (informally) at the differential operator $A$ describing the wave equation (1) by applying the classical substitutions $x_{1}=\rho\partial_{t}w$ and $x_{2}=w$ and obtain:
$\partial_{t}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\mathrm{I}\\ -\mathrm{I}&0\end{pmatrix}\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{=\mathrlap{A}}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}+\begin{pmatrix}0&0\\ -a&-b\frac{1}{\rho}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$
@@
$\operatorname{dom}A\coloneqq\left\{\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in X\,\middle|\,\mathopen{}\tfrac{1}{\rho}x_{2}\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\ \text{and}\ T\nabla x_{1}\in\mathrm{H}(\operatorname{div},\Omega)\right\}\text{.}$
@@
Let $(\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix})_{n\in\mathbb{N}}$ be a sequence in $\operatorname{dom}A$ that converges to $\begin{pmatrix}x\\ y\end{pmatrix}\in X$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{X}$ such that also $(A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix})_{n\in\mathbb{N}}$ converges to $\begin{pmatrix}v\\ w\end{pmatrix}\in X$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{X}$. Note that $A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}=\begin{pmatrix}\frac{1}{\rho}y_{n}\\ \operatorname{div}T\nabla x_{n}\end{pmatrix}$ implies that $\frac{1}{\rho}y_{n}$ converges to $v$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ and therefore $v=\frac{1}{\rho}y$. Since $x_{n}$ converges to $x$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ we conclude that $T\nabla x_{n}$ converges to $T\nabla x$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$. Hence, the closedness of $\operatorname{div}$ implies $w=\operatorname{div}T\nabla x$, which shows the closedness of $A$. ∎
$A$ is what we have called $A_{0}^{\ast}$ in Definition 2.7 for an (at this point still undefined) operator $A_{0}$. [^11] In fact, we will forego the definition of $A_{0}$ entirely. As outlined in Section 2, we want to attach boundary conditions to $A$ by identifying a suitable boundary triple and then define $A_{\Theta}$. For our boundary operators we define:
@@
The ultimate goal of this section is to show that $\bigl{(}\bigl{(}\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\bigr{)},B_{1},B_{2}\bigr{)}$ is a boundary triple for $A$. From this, well-posedness of Equation 1 will easily follow. According to Definition 2.7 we have to verify:
@@
    $B\coloneqq\begin{pmatrix}B_{1}\\ B_{2}\end{pmatrix}$ is onto.
@@
    $A$ and $B$ satisfy an abstract Green identity.
To verify surjectivity of $B$ it is enough to show that the individual $B_{1}$ and $B_{2}$ are surjective, as they act on different components of $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$. We first need a very useful result about solutions of the (generalized) Dirichlet equation with boundary condition provided by $B_{1}$. The result is a slight modification of [22, Thm. 5.5]:
@@
For any $g\in\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ there exists $w\in\{f\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\,|\,\mathopen{}T\nabla f\in\mathrm{H}(\operatorname{div},\Omega)\}$ such that
@@
In particular, $w=\gamma_{0}^{\ast}g$ is given by the adjoint
$\gamma_{0}^{\ast}\colon\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})\to\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\quad\text{of}\quad\gamma_{0}\colon\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\to\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}),$
@@
$\langle f,g\rangle_{\nabla,\gamma_{0}}\coloneqq\langle T\nabla f,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}f,\gamma_{0}g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}.$
@@
Note that $\gamma_{0}\colon\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\to\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})$ is bounded. Thus, its adjoint is defined on all of $\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ and is bounded. Moreover, $\gamma_{0}^{\ast}g\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$ as this is the range of $\gamma_{0}^{\ast}$. By definition of the adjoint we have for $v\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$
@@
If we choose $v\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)$ we obtain
@@
$=\langle\gamma_{\nu}T\nabla\gamma_{0}^{\ast}g+k_{1}\gamma_{0}\gamma_{0}^{\ast}g,\gamma_{0}v\rangle_{\mp\frac{1}{2}},$
which in turn implies $w=\gamma_{0}^{\ast}g$, finishing the proof. ∎
@@
By the definition of the inner product in $X$ we have
@@
Applying integration by parts on the $\operatorname{div}T\nabla$ operators gives (we suppress the index for the inner product in the following and just distinguish between inner products and dual pairings)
@@
$=\langle B_{1}x,B_{2}y\rangle_{\mp\frac{1}{2}}+\langle B_{2}x,B_{1}y\rangle_{\pm\frac{1}{2}}$
To avoid having to define $A_{0}$ (from Definition 2.7) itself and having to verify that $A_{0}^{\ast}=A$, we will simply prove $A^{\ast}\subseteq-A$. This way, $A_{0}\coloneqq-A^{\ast}$ has all necessary properties of $A_{0}$, i.e., skew-symmetry, dense domain and closedness. [^12] Alternatively one can set $A_{0}\coloneqq-A$ with $\operatorname{dom}A_{0}\coloneqq\ker B_{1}\cap\ker B_{2}$, cf. [19, Lem. 2.4.5]. However, this requires careful handling to avoid circular arguments.
@@
$A^{\ast}\subseteq-A$ and $A$ is densely defined
@@
$\langle Ax,y\rangle_{X}=\langle x,z\rangle_{X}$
@@
This holds true for all $x_{2}\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)$, hence $T\nabla y_{1}\in\mathrm{H}(\operatorname{div},\Omega)$ and $z_{2}=-\operatorname{div}T\nabla y_{1}$.
For arbitrary $g\in\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ we let $x_{1}=\gamma_{0}^{\ast}g$. Appealing to Proposition 3.2 we conclude $B_{1}\begin{pmatrix}x_{1}\\ 0\end{pmatrix}=g$ and $\operatorname{div}T\nabla x_{1}=0$. This assures $x=\begin{pmatrix}x_{1}\\ 0\end{pmatrix}\in\operatorname{dom}A$, which allows us to calculate
@@
Since $g\in\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ was arbitrary, we conclude $\gamma_{0}z_{1}=0$. Appealing to the Helmholtz decomposition $\mathrm{L}^{2}(\Omega)=T\nabla\mathring{\mathrm{H}}^{1}(\Omega)\oplus_{T^{-1}}\ker\operatorname{div}$ from Theorem 2.6 we know that $\operatorname{ran}(\operatorname{div}T\mathring{\nabla})=\operatorname{ran}(\operatorname{div})=\mathrm{L}^{2}(\Omega)$. For $x=\begin{pmatrix}x_{1}\\ 0\end{pmatrix}$ for arbitrary $x_{1}\in\{f\in\mathring{\mathrm{H}}^{1}(\Omega)\,|\,\mathopen{}T\nabla x_{1}\in\mathrm{H}(\operatorname{div},\Omega)\}$ we have $x\in\operatorname{dom}A$ and by (9)
@@
Altogether we have shown $y\in\operatorname{dom}A$ and $z=-Ay$, i.e., $A^{\ast}\subseteq-A$
Note that $A^{\ast}$ is not multi-valued as $A^{\ast}\subseteq-A$. This implies $(\operatorname{dom}A)^{\perp}=\{0\}$, cf. [19, Lem. 2.2.8] or equivalently $\overline{\operatorname{dom}A}=X$. ∎
With this, all initially outlined steps for verification of a boundary triple are in place.
@@
By definition of a boundary triple, we have to check, that an abstract Green identity holds (verified in Proposition 3.4) and that $B$ is onto (verified in Corollary 3.3). ∎
The remainder of this article heavily relies on exploiting Theorem 3.6. The first step is verifying well-posedness of the initial problem (1).
@@
For this we simply have to specify the boundary relation $\Theta$ from Equation 6. We set
[…]
Then the collection of boundary conditions from (3) can be written as $B_{1}x+k_{2}B_{2}x=0$ on $\tilde{\Gamma}$. We use the notation $\Theta$ just to fit the usual parameter for boundary triples. In particular
$B_{1}x+k_{2}B_{2}x=0\quad\iff\quad k_{1}w+\nu\cdot T\nabla w+k_{2}\partial_{t}w=0.$
If one replaces $k_{1},k_{2}\in\mathcal{L}(\mathrm{L}^{2}(\tilde{\Gamma}))$ with suitable multiplication operators arising from functions by the same name, the equation to the right contains the second to the fifth condition of (3) in one line by taking into account that $k_{1}$ and $k_{2}$ may vanish on certain parts of $\tilde{\Gamma}$. The condition $w=0$ on $\Gamma_{0}$ is already encoded in the domain of the operator $A$ and completes the set of equations from (3). Hence, the operator that encodes all boundary conditions is $A_{\Theta}$ defined by (6) (with $A_{0}^{\ast}=A$ from (7) and $\Theta=k_{2}$), i.e.,
@@
By Theorem 3.6 $\bigl{(}(\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})),B_{1},B_{2}\bigr{)}$ is a boundary triple for $A$. Hence, Proposition 2.8 and the Lumer–Philips theorem (Theorem 2.5) imply the claim. ∎
Theorem 3.7 assures well-posedness of the wave equation without the perturbation terms $a$ and $b$. Note that we assumed the initial conditions be in the state space $X$ already in Section 1. For the full problem Equation 1 we define the perturbation
$S\colon X\to X,\quad\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\mapsto\begin{pmatrix}0\\ -ax_{1}-\frac{1}{\rho}bx_{2}\end{pmatrix}$
for $a,b\in\mathrm{L}^{\infty}(\Omega)$ we obtain a bounded linear operator. Then Equation 1 becomes
@@
and we have unique solvability appealing to standard perturbation theory (e.g., [9, Thm. 1.3]). One can also allow more general $a,b$ as long as $S$ remains suitably relatively $A_{\Theta}$ bounded, cf. [9, Thm. 2.7] for a suitable perturbation result.

### v2

@@ 3. Well-posedness
In [18] the existence of a boundary triple associated to the wave equation in the Dirac representation was shown. In this section, we show that we can also associate a boundary triple to the wave equation in the Lagrange representation. This will allow us to parameterize all maximally dissipative boundary conditions.
Recall the decomposition of $\partial\Omega$ into $\Gamma_{0}$, $\Gamma_{1}$, $\Gamma_{2}$ $\Gamma_{3}$ and $\Gamma_{4}$. Since the case where $\Gamma_{0}$ is equal to $\partial\Omega$ simply corresponds to the wave equation with Dirichlet boundary, which is well-studied, we exclude that case. In particular, we will see that the boundary conditions (3) represent one of these maximally dissipative boundary conditions. For the following, we only have to distinguish between $\Gamma_{0}$ and the combined other parts of the boundary. Hence, let us abbreviate $\tilde{\Gamma}\coloneq\partial\Omega\setminus\overline{\Gamma_{0}}$, which is essentially the same as $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\cup\Gamma_{4}$.
@@
$X\coloneq\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\times\mathrm{L}^{2}(\Omega).$
We use the equivalent norm for $\mathrm{H}^{1}(\Omega)$ given by the Friedrichs/Poincaré inequality (A.1) and define
$\langle x,y\rangle_{X}\coloneq\big\langle\tfrac{1}{\rho}x_{2},y_{2}\big\rangle_{\mathrm{L}^{2}(\Omega)}+\langle T\nabla x_{1},\nabla y_{1}\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}x_{1},\gamma_{0}y_{1}\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}\text{.}$
Equivalence to the canonical inner product on $X$ is assured by virtue of the Friedrichs/Poincaré inequality and the assumptions on $\rho$, $T$ and $k_{1}$. The inner product on $X$ is the sum of the energy inner product favored in port-Hamiltonian formulations plus the inner product on the boundary, which is necessary to prevent constant functions in the second argument to have vanishing norm. We arrive (informally) at the differential operator $A$ describing the wave equation (1) by applying the classical substitutions $x_{1}=\rho\partial_{t}w$ and $x_{2}=w$ and obtain:
$\partial_{t}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\mathrm{I}\\ -\mathrm{I}&0\end{pmatrix}\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{=\mathrlap{A}}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}+\begin{pmatrix}0&0\\ -a&-b\frac{1}{\rho}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$.
@@
$\operatorname{dom}A\coloneq\left\{\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in X\,\middle|\,\mathopen{}\tfrac{1}{\rho}x_{2}\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\ \text{and}\ T\nabla x_{1}\in\mathrm{H}(\operatorname{div},\Omega)\right\}\text{.}$
@@
Let $(\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix})_{n\in\mathbb{N}}$ be a sequence in $\operatorname{dom}A$ that converges to $\begin{pmatrix}x\\ y\end{pmatrix}\in X$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{X}$ such that also $(A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix})_{n\in\mathbb{N}}$ converges to $\begin{pmatrix}v\\ w\end{pmatrix}\in X$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{X}$. Note that $A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}=\begin{pmatrix}\frac{1}{\rho}y_{n}\\ \operatorname{div}T\nabla x_{n}\end{pmatrix}$ implies that $(\frac{1}{\rho}y_{n})_{n}$ converges to $v$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ and therefore $v=\frac{1}{\rho}y$. Since $(x_{n})_{n}$ converges to $x$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$, we conclude that $(T\nabla x_{n})_{n}$ converges to $T\nabla x$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$. Hence, the closedness of $\operatorname{div}$ implies $w=\operatorname{div}T\nabla x$, which shows the closedness of $A$. ∎
AA is what we have called $A_{0}^{\ast}$ in Definition 2.7 for an (at this point still undefined) operator $A_{0}$. [^12] In fact, we will pretermit the definition of $A_{0}$ entirely. As outlined in Section 2, we want to attach boundary conditions to $A$ by identifying a suitable boundary triple and then define $A_{\Theta}$. For our boundary operators we define:
@@
The ultimate goal of this section is to show that $\bigl(\bigl(\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\bigr),B_{1},B_{2}\bigr)$ is a boundary triple for $A$. This will allow us to attach the boundary conditions (3) to $A$ via a relation $\Theta$ essentially given by $k_{2}$ as described in Section 2.5. From there, well-posedness of problem (1) will easily follow. According to Definition 2.7, for a boundary triple we have to verify:
@@
    $B\coloneq\begin{pmatrix}B_{1}\\ B_{2}\end{pmatrix}$ is onto.
@@
    AA and $B$ satisfy an abstract Green identity.
To verify surjectivity of $B$ it suffices to show that the individual $B_{1}$ and $B_{2}$ are surjective, since they act on different components of $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$. We first establish a useful result about solutions of the (generalized) Dirichlet problem with boundary condition provided by $B_{1}$. The following proposition is a slight modification of [27, Thm. 5.5]:
@@
[…]
@@
In particular, $w\coloneq\gamma_{0}^{\ast}g$ given by the adjoint
[…]
@@
$\langle f,g\rangle_{\nabla,\gamma_{0}}\coloneq\langle T\nabla f,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}f,\gamma_{0}g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}$,
is a solution.
@@
[…]
@@
If we choose $v\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)$, we obtain
@@
$=\langle\gamma_{\nu}T\nabla\gamma_{0}^{\ast}g+k_{1}\gamma_{0}\gamma_{0}^{\ast}g,\gamma_{0}v\rangle_{\mp\frac{1}{2}}.$
Hence $w\coloneq\gamma_{0}^{\ast}g$ is a solution. ∎
@@
By the definition of the inner product in $X$ we have:
@@
Applying integration by parts on the $\operatorname{div}T\nabla$ operators gives (we suppress the index for the inner product in the following and just distinguish between inner products and dual pairings):
@@
$=\langle B_{1}x,B_{2}y\rangle_{\mp\frac{1}{2}}+\langle B_{2}x,B_{1}y\rangle_{\pm\frac{1}{2}}.$
To avoid having to define $A_{0}$ (from Definition 2.7) itself and having to verify that $A_{0}^{\ast}=A$, we will simply prove $A^{\ast}\subseteq-A$. This way, $A_{0}\coloneq-A^{\ast}$ has all necessary properties of $A_{0}$, i.e., skew-symmetry, dense domain and closedness. [^13] Alternatively, one can set $A_{0}\coloneq-A$ with $\operatorname{dom}A_{0}\coloneq\ker B_{1}\cap\ker B_{2}$, cf. [24, Lem. 2.4.5]. However, this requires careful handling to avoid circular arguments.
@@
$A^{\ast}\subseteq-A$ and $A$ is densely defined.
@@
$\langle Ax,y\rangle_{X}=\langle x,z\rangle_{X},$
@@
This holds true for all $x_{2}\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)$; hence $T\nabla y_{1}\in\mathrm{H}(\operatorname{div},\Omega)$ and $z_{2}=-\operatorname{div}T\nabla y_{1}$.
[…]
@@
[…]
@@
Altogether we have shown $y\in\operatorname{dom}A$ and z=Ayz=-Ay, i.e., $A^{\ast}\subseteq-A$.
Note that $A^{\ast}$ is not multi-valued, since $A^{\ast}\subseteq-A$. This implies $(\operatorname{dom}A)^{\perp}=\{0\}$, cf. [24, Lem. 2.2.8], or equivalently $\overline{\operatorname{dom}A}=X$. ∎
With this, all initially outlined steps for the verification of a boundary triple are in place.
@@
By definition of a boundary triple, we have to check that an abstract Green identity holds (verified in Proposition 3.4) and that $B$ is onto (verified in Corollary 3.3). ∎
The remainder of this article heavily relies on exploiting Theorem 3.6. The first step is to establish well-posedness of the initial problem (1).
@@
For this we simply have to specify the boundary relation $\Theta$ from (6). We set
$\Theta\coloneq k_{2}\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\tilde{\Gamma}))\ \text{(as an operator)}.$
Then the collection of boundary conditions from (3) can be written as $B_{1}x+k_{2}B_{2}x=0$ on $\tilde{\Gamma}$. In particular,
$B_{1}x+k_{2}B_{2}x=0\quad\iff\quad k_{1}w+\nu\cdot T\nabla w+k_{2}\partial_{t}w=0.$ (10)
If one replaces $)k_{1},k_{2}\in\mathcal{L}(\mathrm{L}^{2}(\tilde{\Gamma}$)) with suitable multiplication operators arising from functions by the same name, the equation to the right contains the second to the fifth condition of (3) in one line by taking into account that $k_{1}$ and $k_{2}$ may vanish on certain parts of $\tilde{\Gamma}$. The condition $w=0$ on $\Gamma_{0}$ is already encoded in the domain of the operator $A$ and completes the set of equations from (3). Hence, the operator that encodes all boundary conditions is $A_{\Theta}$ defined by (6) (with $A_{0}^{\ast}=A$ from (7) and $\Theta=k_{2}$), i.e.,
@@
By Theorem 3.6 $\bigl((\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})),B_{1},B_{2}\bigr)$ is a boundary triple for $A$. Hence, Proposition 2.8 and the Lumer–Philips Theorem 2.5 imply the claim. ∎
Theorem 3.7 assures well-posedness of the wave equation without the perturbation terms $a$ and $b$. Note that we assumed the initial conditions be in the state space $X$ already in Section 1. For the full problem (1) we define the perturbation
$S\colon X\to X,\quad\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\mapsto\begin{pmatrix}0\\ -ax_{1}-\frac{1}{\rho}bx_{2}\end{pmatrix}$.
For $a,b\in\mathrm{L}^{\infty}(\Omega)$ we obtain a bounded linear operator. Then the differential equation from (1) becomes
@@
and we have unique solvability appealing to standard perturbation theory (e.g., [12, Thm. 1.3]). One can also allow more general $a,b$ as long as $S$ remains suitably relatively $A_{\Theta}$ bounded, cf. [12, Thm. 2.7] for a suitable perturbation result.